OpenStudy (anonymous):

why isnt this working? A water truck's tank has internal dimensions of 11 meters in length and 2 meters in diameter. If the truck's load of pure water is 70 percent of full capacity, how heavy is the load in pounds? Water's density is 62.4 lb/ft^3. In kilograms, what is the weight of the load if the truck carries liquid nitrogen instead of water? The liquid nitrogen's density is 0.808 g/ml. I found the first part...but isnt the second part (pi)(1^2)(11)(.7)(35.3)(808)

OpenStudy (asnaseer):

where did the 35.3 come from in (pi)(1^2)(11)(.7)(35.3)(808)?

OpenStudy (anonymous):

1m^3 = 35.3ft^3

OpenStudy (asnaseer):

but why are you converting it to those units?

OpenStudy (anonymous):

to get to kilograms

OpenStudy (asnaseer):

its given you the density as 0.808 g/ml - whereas the first part was in ib/ft^3

OpenStudy (asnaseer):

so all you need to do is find out how many ml there are in each cubic meter, and convert grams to kilograms

OpenStudy (asnaseer):

1 ml = \(1 cm^3\)

OpenStudy (anonymous):

is the volume formula useless too

OpenStudy (asnaseer):

no - that is correct

OpenStudy (asnaseer):

The liquid nitrogen's density is 0.808 g/ml = 0.808 g/cm^3 Now just convert the grams into kilograms by dividing by 1000

OpenStudy (anonymous):

so my final set up should be (pi)(1^2)(11)(.7)(35.3)(808)/1000?

OpenStudy (asnaseer):

no - remove the 35.3

OpenStudy (asnaseer):

Volume of full tank = \(\pi*1^2*11\) So 70% capacity will have a volume of \(\pi*1^2*11*.7\) - so you have the volume correct

OpenStudy (asnaseer):

the question is a little tricky because they gave you the density of water in ib/ft^3 whereas the density of nitrogen is given in g/ml.

OpenStudy (asnaseer):

so you have to be careful with the units.